OpenStudy (anonymous):

help solve: e^x(e^x - 5) = -4

OpenStudy (anonymous):

\[e^x(e^x - 5) = -4\] and then I went to: \[(e^x)^2 - 5(e^x) = -4\] but I feel like I'm on the wrong path. Please help!

OpenStudy (anonymous):

you're on the right track. next step:\[e^{2x} - 5e^x + 4= 0\]and like a polynomial, you can factor it:\[(e^x - 4)(e^x - 1) = 0\] zero product rule: 0 when e^x = 4 or e^x = 1 x = ln(4) or ln(1)

OpenStudy (anonymous):

thanks! how did you get x = ln4 or ln1 though?

OpenStudy (anonymous):

\[e^x = 4\]\[\ln(e^x) = \ln(4)\] and \[\ln(e^x) = x\] LN is the inverse operation of the exponential function. you can also play around and test with your calculator to confirm this

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

glad i could help :) let me know if you have follow-up questions :)